### v1

@@ 3.2 Unavoidable loss due to errors in allocation schedule
Motivated by this example, our key insight is that, for any static allocation $X$, there exists a worst-case loss due to errors in $\sigma$, denoted by $\overline{\Delta}(X)$, that the algorithm incurs. As a result, rather than having a budget of $B$ items, the algorithm has an effective budget of $B-\overline{\Delta}(X)$ items. Under this effective budget, any feasible stationary allocation must set $X$ such that $\overline{N}X\leq{B-\overline{\Delta}(X)}$, where $\overline{N}$ is a high-probability upper bound on $N$. Noting that $X=0$ is always a feasible solution to this inequality, a natural choice is to set:
$\underline{X}=\sup\mathopen{}\mathclose{{}\left\{X\ \mid\ X\leq\frac{B-\overline{\Delta}(X)}{\overline{N}}}\right\},$ (4)
@@
if this supremum is achieved. When $\overline{\Delta}(X)=0$ for all $X$ (i.e., either items don’t perish or predictions are perfect), $\underline{X}=B/\overline{N}$, and we recover the conservative allocation of the no-perishing setting (Sinclair et al., , 2022). The notion of $\sigma$-induced loss plays a central role in our results.
Definition 3.3 ($\sigma$-induced loss).
We define $\mathcal{L}^{\textsf{perish}}=\frac{B}{\overline{N}}-\underline{X}$ to be any algorithm’s $\sigma$-induced loss. We moreover term $T\mathcal{L}^{\textsf{perish}}$ to be the cumulative $\sigma$-induced loss.
@@
By (4), given the worst-case perishing loss $\overline{\Delta}(X)$ for any allocation $X$, one can compute $\underline{X}$ via line search. Obtaining tight bounds on this quantity, however, is the challenging piece. To see this, note that for any algorithm, the quantity of goods that perished by the end of the horizon is:
@@
where $\mathcal{B}_{t}^{alg}$ denotes the set of remaining items at the beginning of period $t$, $B_{t}^{alg}(b)$ is the quantity of item $b$ remaining at the beginning of period $t$, and $X_{t}^{alg}(b)$ is the quantity of item $b$ given out in period $t$. Since $X_{t}^{alg}(b)$ depends on the perishing realizations of previous rounds, computing this quantity requires the ability to simulate sufficiently many replications of the static allocation process under $X$, for all $X\in[0,B/\overline{N}]$, and for each of these replications to compute the number of unallocated goods that perished by the end of the horizon under this allocation, an approach which fails to scale.
To tackle this difficulty, it will be useful for us to consider a “slow” consumption process, in which $\underline{N}_{\leq t}$ — a high-probability lower bound on $N_{\leq t}$ — individuals arrive before $t+1$, $\underline{N}_{\leq t}{X}$ items are allocated up to period $t\in[T]$, and no items perish. For $b\in\mathcal{B}$, we let ${\tau}_{b}(1\mid X,\sigma)$ be the period in which $b$ would have been entirely allocated under this slow consumption process. Formally,
${\tau}_{b}(1\mid X,\sigma)=\inf\mathopen{}\mathclose{{}\left\{t\geq 1:\underline{N}_{\leq t}X\geq\sigma(b)}\right\}.$ (6)
${\tau}_{b}(1\mid X,\ \sigma)$ represents an upper bound on the time an algorithm using static allocation $X$ would allocate $b$, since items ranked higher than $b$ may have perished, thus decreasing the time at which $b$ is allocated. We define $\mu(X)=\sum_{b\in\mathcal{B}}\mathbb{P}\mathopen{}\mathclose{{}\left(T_{b}<\min\mathopen{}\mathclose{{}\left\{T,{\tau}_{b}(1\mid X,\sigma)}\right\}}\right)$ and let
$\overline{\Delta}({X})=\min\mathopen{}\mathclose{{}\left\{B,\mu(X)+\textsc{Conf}^{P}_{1}(\mu(X))}\right\},$ (7)
@@
Figure 2: Illustrating the $\underline{X}$ construction (4) for the toy instance in Example 3.6. The dashed line corresponds to the line $Y=X$, and the solid line to $(B-\overline{\Delta}(X))/\overline{N}$. Here, $\underline{X}$ is represented by the red star, the point at which the solid and dashed lines intersect.
Remark 3.4.
We henceforth assume for simplicity that the supremum on the right-hand side of (4) is attained. This is without loss to our results, since our bounds depend on the $\sigma$-induced loss $\mathcal{L}^{\textsf{perish}}$. Hence, if the supremum fails to be attained, one can define $\underline{X}=X^{*}-\epsilon$, where $X^{*}$ is the point of discontinuity and $\epsilon=o(1)$ guarantees feasibility of $\underline{X}$.
Remark 3.5.
Example 3.6 illustrates the $\underline{X}$ construction for a toy instance.
Example 3.6.
Consider a setting where $B=T=4$, $|\Theta|=1$, and $N_{t}=1$ for all $t\in[T]$, with the following perishing time distributions for each item:
$T_{1}=\begin{cases}1\quad\text{w.p. }1/2\\ 2\quad\text{w.p. }1/2\end{cases}\quad T_{2}=\begin{cases}1\quad\text{w.p. }1/2\\ 4\quad\text{w.p. }1/2\end{cases}\quad T_{3}=\begin{cases}2\quad\text{w.p. }1/2\\ 3\quad\text{w.p. }1/2\end{cases}\quad T_{4}=\begin{cases}3\quad\text{w.p. }1/2\\ 4\quad\text{w.p. }1/2\end{cases}.$
Let $\sigma(b)=b$ for all $b$ (i.e., items are allocated in increasing order of expected perishing time, with ties broken in favor of earliest possible perishing time). Fig. 2 illustrates the solution to (4) for this instance, with $\textsc{Conf}_{1}^{P}(\mu(X))=0$ for all $X$. Observe that $\underline{X}=0.25$, whereas the proportional allocation $B/N=1$.
Observe that the perishing process described in Example 3.6 is not almost surely offset-expiring, since items 1 and 2 both perish at the end of period 1 with probability 1/4. However, lack of offset-expiry is not the reason that any online algorithm incurs an additional perishing-related loss. Theorem 3.7 establishes that any online algorithm’s performance necessarily scales with the $\sigma$-induced loss, even under offset-expiry.
Theorem 3.7.
There exists an offset-expiring instance such that, for any online algorithm that is feasible with probability at least $\alpha$, the following holds with probability at least $\alpha$:
@@
Proof.
Consider an instance with $B=T$, $|\Theta|=1$ and $N_{t}=1$ for all $t$. Suppose moreover that resources have deterministic perishing times, with $T_{b}=b$ for all $b$, and $\sigma(b)=T+1-b$. ($T_{b}=b$ implies that the process is offset-expiring.) Since the perishing and demand processes are deterministic, we let $\textsc{Conf}_{1}^{P}(\mu(X))=0$ for all $X$, and $\overline{N}=N$. For ease of notation, we omit the dependence of all quantities on $\theta$ in the remainder of the proof. The following lemma states that under this flipped ordering, any online algorithm is severely limited in its total allocation.
Lemma 3.8.
@@
Proof.
For any feasible algorithm, there must exist an available unit in period $T$. Since the only unit that has not perished by $t=T$ is $b=1$, it must be that this unit is available in period $T$. Thus, we must have $\sum_{t}X_{t}^{alg}\leq 1$ (else, $b=1$ will have been allocated before $T$). ∎
Lemma 3.8 implies that any feasible stationary algorithm, which allocates a fixed amount $X_{t}^{alg}=X$ for all $t$, must have $X\leq\frac{1}{N}$. We use this fact to bound $\overline{\Delta}(X)$, for any feasible stationary allocation.
Lemma 3.9.
@@
$\overline{\Delta}(X)=\sum_{b}\mathds{1}\mathopen{}\mathclose{{}\left\{T_{b}<\min\{T,{\tau}_{b}(1\mid X,\sigma)\}}\right\}\geq\sum_{b}\mathds{1}\mathopen{}\mathclose{{}\left\{b<\min\{T,T({T+1-b})\}}\right\}=\sum_{b}\mathds{1}\mathopen{}\mathclose{{}\left\{b<T}\right\}$
@@
where the inequality uses the lower bound on ${\tau}_{b}(1\mid X,\sigma)$ in addition to the assumption that $T_{b}=b$. ∎
Putting Lemma 3.8 and Lemma 3.9 together, we have:
$\underline{X}:=\sup\{X:X\leq\frac{B-\overline{\Delta}(X)}{N}\}\leq\sup\{X:X\leq\frac{T-(T-1)}{T}\}=\frac{1}{T}\\ \implies\mathcal{L}^{\textsf{perish}}\geq 1-1/T.$
We now show the lower bounds on $\Delta_{\text{\it EF}}$ and $\Delta_{\text{\it efficiency}}$. By Lemma 3.8, $\sum_{t}X_{t}^{alg}\leq 1$, which implies that $\min_{t}X_{t}^{alg}\leq\frac{1}{T}$. Hence, $\Delta_{\text{\it EF}}=\max_{t}|1-X_{t}^{alg}|\geq 1-\frac{1}{T}=\mathcal{L}^{\textsf{perish}}$. Moreover, $\Delta_{\text{\it efficiency}}=T-\sum_{t}X_{t}^{alg}\geq T-1=T\mathcal{L}^{\textsf{perish}}$. ∎
4 The Algorithm
Our algorithm, Perishing-Guardrail, takes as input $(i)$ a desired bound on envy $L_{T}$, and $(ii)$ a high-probability parameter $\delta$. The algorithmic approach is tackled in three steps:
@@ 4 The Algorithm
    Constructing a static allocation (also referred to as baseline allocation or lower guardrail), $\underline{X}$, under which the algorithm doesn’t run out of budget with high probability. Motivated by Section 3.2, we let
    $\underline{X}=\sup\mathopen{}\mathclose{{}\left\{X\mid X\leq\frac{B-\overline{\Delta}(X)}{\overline{N}}}\right\},$
    with $\overline{N}=\mathbb{E}[N]+\textsc{Conf}_{0,T}^{N}$, for an appropriately defined high-probability confidence term $\textsc{Conf}_{0,T}^{N}$.
@@
    Determining an appropriate threshold condition that indicates when to allocate $\overline{X}$.
Though the above approach is similar to the guardrail algorithm proposed by Sinclair et al., (2022) for the no-perishing setting, we emphasize that identifying the appropriate static allocation and threshold condition under perishing uncertainty poses significant challenges that do not exist in the classical setting. In this latter setting, the natural static allocation that guarantees budget-feasibility is the proportional allocation in a high-demand regime, i.e., $\underline{X}=B/\overline{N}$. Part of the reason this is easily handled is the fact that arrival uncertainty is exogenous, i.e. it is invariant to the decisions made by the algorithm. On the other hand, uncertainty around perishing is endogenous: as discussed in Section 3, though the distribution around perishing times is fixed, how many — and which — items perish depends heavily on the rate at which items are being allocated, which itself depends on the rate at which items perish. The threshold condition we next describe must contend with this knife-edge effect.
Determining the threshold condition.
Recall, the “aggressive” allocation $\overline{X}$ will be used to improve our algorithm’s efficiency, at the cost of higher envy. In each period, our algorithm checks whether there is enough budget remaining to accommodate $(i)$ an aggressive allocation in the current period, $(ii)$ a conservative allocation in all future periods, under high demand, and $(iii)$ perishing that may occur under future conservative allocations. The main challenge here lies in estimating $(iii)$: over-optimism runs an increased risk of exhausting the budget early, due to the same phenomenon as that described in Section 3.2, whereas over-pessimism fails to take advantage of efficiency gains from aggressively allocating.
For $t\in[T]$, we let $\overline{P}_{t}$ denote our algorithm’s worst-case perishing prediction. As above, the “slow” consumption process will allow us to obtain a closed-form characterization of $\overline{P}_{t}$. In particular, for $t\in[T]$, $b\in\mathcal{B}$, we consider the notion of “latest allocation time” after $t$:
${\tau}_{b}(t\mid\underline{X},\sigma)=\inf\mathopen{}\mathclose{{}\left\{t^{\prime}\geq t:\underline{N}_{<t}\underline{X}+\underline{N}_{[t,t^{\prime}]}\underline{X}\geq\sigma(b)}\right\},$ (8)
where $\underline{N}_{[t,t^{\prime}]}=\mathbb{E}\mathopen{}\mathclose{{}\left[N_{[t,t^{\prime}]}}\right]-\textsc{Conf}_{t,t^{\prime}}^{N}$ for an appropriately defined high-probability confidence term $\textsc{Conf}_{t,t^{\prime}}^{N}$. In words, $\underline{N}_{<t}\underline{X}+\underline{N}_{[t,t^{\prime}]}\underline{X}$ corresponds to the least amount the algorithm could have consumed by $t^{\prime}$ (either by allocating or via perishing). Hence, if $b$ is in the set of remaining items at the beginning of period $t$, with high probability it will be allocated before ${\tau}_{b}(t\mid\underline{X},\sigma)$. Via similar logic, we let ¯$\overline{\mathcal{B}}_{t}=\{\sigma^{-1}(\lceil\underline{N}_{<t}\underline{X}\rceil),\ldots,\sigma^{-1}(B)\}$ be the set of items remaining under this slow process, and define the expected number of items that perish from $t$ onwards as:
$\eta_{t}=\sum_{b\in\overline{\mathcal{B}}_{t}}\mathbb{P}\mathopen{}\mathclose{{}\left(t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\}}\right).$ (9)
The pessimistic forecast of future spoilage is then defined as:
$\begin{cases}\overline{P}_{t}&=\min\mathopen{}\mathclose{{}\left\{\overline{P}_{t-1},\eta_{t}+\textsc{Conf}_{t}^{P}(\eta_{t})}\right\}\quad\forall\ t\in[T],\\ \overline{P}_{0}&=B,\end{cases}$ (10)
@@
for an appropriately defined confidence term $\textsc{Conf}_{t}^{P}(\cdot)$. Note that $\overline{P}_{1}=\overline{\Delta}(\underline{X})$ (see Eq. 7).
@@
We present our algorithm, Perishing-Guardrail, in Algorithm 1. For $t\in[T]$, $\mathcal{B}^{alg}_{t}$ is used to denote the set of remaining resources at the beginning of time $t$, and $B_{t}^{alg}$ the quantity of remaining resources at the beginning of the period. Moreover, let $\mathcal{A}_{t}$ be the set of items allocated in round $t$, and $\textsc{PUA}_{t}^{alg}$ the quantity of unallocated items that perished at the end of round $t$.
@@
Input: Budget $B=B_{1}^{alg}$, allocation schedule $\sigma$, envy parameter $L_{T}$, arrival confidence terms $(\textsc{Conf}_{t,t^{\prime}}^{N})_{t,t^{\prime}\in\{0,\ldots,T\}}$, perishing confidence terms $(\textsc{Conf}_{t}^{P}(\cdot))_{t\in\{1,\ldots,T\}}$, and perishing inputs $(\eta_{t})_{t\in[T]}$ given by (9)
@@
Compute $\underline{X}=\sup\mathopen{}\mathclose{{}\left\{X\mid X\leq\frac{B-\overline{\Delta}(X)}{\overline{N}}}\right\}$ and set $\overline{X}=\underline{X}+L_{T}$.
@@
       Compute $\overline{P}_{t}=\min\mathopen{}\mathclose{{}\left\{\overline{P}_{t-1},\eta_{t}+\textsc{Conf}_{t}^{P}(\eta_{t})}\right\}$ // Compute ‘‘worst-case’’ future perishing
@@
       else if $B_{t}^{alg}-N_{t}\overline{X}\geq\underline{X}(\mathbb{E}\mathopen{}\mathclose{{}\left[N_{>t}}\right]+\textsc{Conf}_{t,T}^{N})+\overline{P}_{t}$ then// use upper guardrail

### v2

@@ 3.2 Unavoidable Loss Due to Perishing
Motivated by this example, our key insight is that, for any static allocation $X$, there exists a high-probability worst-case loss, denoted by $\overline{\Delta}(X)$, that any algorithm incurs, due to both errors in $\sigma$, as well as the inherent aggressiveness of the perishing process. As a result, rather than having a budget of $B$ items, the algorithm has an effective budget of $B-\overline{\Delta}(X)$ items. Under this effective budget, any feasible static allocation must set $X$ such that $\overline{N}X\leq{B-\overline{\Delta}(X)}$, where $\overline{N}$ is a high-probability upper bound on $N$. Noting that $X=0$ is always a feasible solution to this inequality, a natural choice is to set:
$\underline{X}=\sup\left\{X\ \mid\ X\leq\frac{B-\overline{\Delta}(X)}{\overline{N}}\right\},$ (9)
@@
if this supremum is achieved. When $\overline{\Delta}(X)=0$ for all $X$ (e.g., either items don’t perish or the process is offset-expiring and predictions are perfect), $\underline{X}=B/\overline{N}$, and we recover the conservative allocation of the no-perishing setting (Sinclair et al. 2023).
@@
By Equation 9, given the worst-case perishing loss $\overline{\Delta}(X)$ for any allocation $X$, one can compute $\underline{X}$ via line search. Obtaining tight bounds on $\overline{\Delta}(X)$, however, is the challenging piece. To see this, note that for any algorithm, the quantity of goods that perished by the end of the horizon is:
@@
where $B_{t}^{alg}(b)$ is the quantity of item $b$ remaining at the beginning of period $t$, and $X_{t}^{alg}(b)$ is the quantity of item $b$ given out in period $t$. Since $X_{t}^{alg}(b)$ depends on the perishing realizations of previous rounds, computing this quantity requires the ability to simulate sufficiently many replications of the static allocation process under $X$, for all $X\in[0,B/\overline{N}]$, and for each of these replications to compute the number of unallocated goods that perished by the end of the horizon under this allocation, an approach which fails to scale.
To tackle this difficulty, it will be useful for us to consider a “slow” consumption process; at a high level, this process will allow us to bound the latest possible time any remaining unit $b$ could be allocated, with high probability. Under this process, $\underline{N}_{\leq t}$ — a high-probability lower bound on $N_{\leq t}$ — individuals arrive before $t+1$, $\underline{N}_{\leq t}{X}$ items are allocated up to period $t\in[T]$, and no items perish. For $b\in\mathcal{B}$, we let ${\tau}_{b}(1\mid X,\sigma)$ be the period in which $b$ would have been entirely allocated under this slow consumption process. Formally,
${\tau}_{b}(1\mid X,\sigma)=\inf\left\{t\geq 1:\underline{N}_{\leq t}X\geq\sigma(b)\right\}.$ (11)
${\tau}_{b}(1\mid X,\ \sigma)$ represents an upper bound on the time an algorithm using static allocation $X$ would allocate $b$, since items ranked higher than $b$ may have perished, thus decreasing the time at which $b$ is allocated. We provide an example of the ${\tau}_{b}(1\mid X,\sigma)$ construction below.
Example 3.5 (continued from Examples 2.1, 3.2 and 3.4)
Fix $X=3/4$. Under deterministic demand, $\underline{N}_{\leq 1}=1$, $\underline{N}_{\leq 2}=3$, and $\underline{N}_{\leq 3}=4$. Then, the latest possible allocation times are given by:
[…]
In words, under the slow consumption process, the latest possible allocation time for the entirety of unit c is $t=2$, since only $0.75$ units have been consumed in the first round. At $t=2$, $2.25$ units have been consumed, meaning that c and a have both been consumed by the end of the second round. Finally, at $t=3$, 33 units have been consumed, meaning that b must also have been consumed by the end of the third round.
Note that a necessary condition for any unit $b$ to have spoiled under static allocation $X$ is for its perishing time $T_{b}$ to be such that $T_{b}<\min\{T,\tau_{b}(1\mid{X},\sigma)\}$. This follows from the fact that, if $T_{b}$ exceeds the latest possible allocation time $\tau_{b}(1\mid{X},\sigma)$, then $b$ must have been allocated before it perished, by definition. Accordingly, we define $\mu(X)=\sum_{b\in\mathcal{B}}\mathbb{P}\left(T_{b}<\min\left\{T,{\tau}_{b}(1\mid X,\sigma)\right\}\right)$ to be an upper bound on the expected number of units that spoil under $X$, and use this quantity to define the worst-case loss $\overline{\Delta}(X)$. Formally:
$\overline{\Delta}({X})=\min\left\{B,\mu(X)+\textsc{Conf}^{P}_{1}(\mu(X))\right\},$ (12)
@@
Figure 2: Illustrating the $\underline{X}$ construction for Example 3.6. The dashed line corresponds to the line $Y=X$, and the solid line to $(B-\overline{\Delta}(X))/\overline{N}$. Here, $\underline{X}$ is represented by the red star, the point at which the solid and dashed lines intersect.
With this machinery in hand, we illustrate the $\underline{X}$ construction for our running example.
Example 3.6 (Continued from Examples 2.1, 3.2, 3.4 and 3.5)
Suppose the perishing distributions are as in Example 3.4, $\overline{N}=4$, and $\textsc{Conf}_{1}^{P}(\mu(X))=0$ for all $X$. Let $X=B/\overline{N}=3/4$. Continuing on from Example 3.5, we have:
$\mu(X)=\mathbb{P}\left(T_{\texttt{a}}<2\right)+\mathbb{P}\left(T_{\texttt{b}}<3\right)+\mathbb{P}\left(T_{\texttt{c}}<2\right)=3/2\implies\overline{\Delta}(X)=3/2.$
Then, $\frac{B-\overline{\Delta}(X)}{\overline{N}}=\frac{3-3/2}{4}=3/8<X$, meaning that setting an allocation of $X=3/4$ is infeasible in this example. In this case, the solution to Equation 9 is $\underline{X}=0.375$, as illustrated in Figure 2.
The above example highlights that perishability forces the decision-maker to allocate less than the asymptotically optimal perishing-free allocation of $B/\overline{N}$, in order to guarantee envy-freeness. The notion of unavoidable perishing-induced loss, which we next define, uses the $\underline{X}$ construction to quantify how much less the decision-maker must allocate to account for spoilage that is out of her control.
Definition 3.7 (Perishing-induced loss)
We define $\mathcal{L}^{\textsf{perish}}=\frac{B}{\overline{N}}-\underline{X}$ to be any algorithm’s perishing-induced loss. We moreover term $T\mathcal{L}^{\textsf{perish}}$ to be the cumulative perishing-induced loss.
Theorem 3.8 establishes that any online algorithm’s performance necessarily scales with $\mathcal{L}^{\textsf{perish}}$.
Theorem 3.8
There exists an instance such that, for any online algorithm that is feasible with probability at least $\alpha$, the following holds with probability at least $\alpha$:
@@
Proof 3.9
Proof. Consider an instance with $B=T$, $|\Theta|=1$ and $N_{t}=1$ for all $t$. Suppose moreover that resources have deterministic perishing times, with $T_{b}=b$ for all $b$, and $\sigma(b)=T+1-b$. Since the perishing and demand processes are deterministic, we let $\textsc{Conf}_{1}^{P}(\mu(X))=0$ for all $X$, and $\overline{N}=N$. For ease of notation, we omit the dependence of all quantities on $\theta$ in the remainder of the proof. The following lemma states that under this flipped ordering, any online algorithm is severely limited in its total allocation.
Lemma 3.10
@@
Proof 3.11
[…]
Lemma 3.10 implies that any feasible static algorithm, which allocates a fixed amount $X_{t}^{alg}=X$ for all $t$, must have $X\leq\frac{1}{N}$. We use this fact to bound $\overline{\Delta}(X)$, for any feasible static allocation.
Lemma 3.12
@@
$\overline{\Delta}(X)=\sum_{b}\mathds{1}\left\{T_{b}<\min\{T,{\tau}_{b}(1\mid X,\sigma)\}\right\}\geq\sum_{b}\mathds{1}\left\{b<\min\{T,T({T+1-b})\}\right\}=\sum_{b}\mathds{1}\left\{b<T\right\}$
@@
[…]
Putting Lemmas 3.10 and 3.12 together, we have:
$\underline{X}:=\sup\{X\mid X\leq\frac{B-\overline{\Delta}(X)}{N}\}\leq\sup\{X\mid X\leq\frac{T-(T-1)}{T}\}=\frac{1}{T}\\ \implies\mathcal{L}^{\textsf{perish}}=1-1/T,$
where the final equality follows from the fact that $\overline{\Delta}(X)=T-1$ for $X=1/T$, and therefore $X\leq\frac{B-\overline{\Delta}(X)}{N}$ is tight at $X=1/T$.
[…]
We conclude the section by noting the implicit dependence of $\mathcal{L}^{\textsf{perish}}$ on the ordering $\sigma$. While $\sigma$ is an input to our model, Theorem 3.8 implies that, if the decision-maker were to optimize over the allocation ordering, the correct choice would be to set $\sigma^{*}\in\arg\min_{\sigma}\mathcal{L}^{\textsf{perish}}(\sigma),$ where we emphasize the dependence of $\mathcal{L}^{\textsf{perish}}$ on the order $\sigma$. Computing such a $\sigma^{*}$, however, is infeasible given the space of $N!$ orderings, as discussed in Section 2.3. In Section 4.4 we identify sufficient conditions on the perishing process and allocation order that guarantee that $\mathcal{L}^{\textsf{perish}}$ vanish asymptotically. We then use these conditions to identify practical and easily interpretable allocation schedules in our numerical experiments (see Section 5.4).
Remark 3.14
We henceforth assume for simplicity that the supremum on the right-hand side of Equation 9 is attained. This is without loss to our results, since our bounds depend on the perishing-induced loss $\mathcal{L}^{\textsf{perish}}$. Hence, if the supremum fails to be attained, one can define $\underline{X}=X^{*}-\epsilon$, where $X^{*}$ is the point of discontinuity and $\epsilon=o(1)$ guarantees feasibility of $\underline{X}$.
Remark 3.15
4 Achieving the Optimal Trade-Off Via Perishing-Aware Guardrails
In this section we design an algorithm that achieves the lower bounds described above. We moreover instantiate our bounds for important special cases.
4.1 Algorithm Description
Our algorithm, Perishing-Guardrail, builds on the guardrail algorithm proposed by Sinclair et al. (2023) for the no-perishing setting. In particular, our algorithm always allocates the same amount to all individuals in a given round (independent of their type). As a result, the only source of envy is due to differences in allocations across rounds, as opposed to across individuals within the same round. Our algorithm controls this source of envy via an input parameter $L_{T}=o(1)$; it then tackles the problem of how much to allocate in each round in three main steps:
@@ 4.1 Algorithm Description
    Constructing a static allocation (also referred to as baseline allocation or lower guardrail), $\underline{X}\in\mathbb{R}_{+}$.
@@
    Determining an appropriate threshold condition that indicates when to allocate aggressively (i.e., in which periods to allocate $\overline{X}$ instead of $\underline{X}$).
At a high level, the lower guardrail $\underline{X}$ is constructed to be the largest feasible allocation for which the hindsight envy is zero (i.e., all individuals receive the same amount in each period), with high probability. For this to hold, it must be that the algorithm never runs out of budget under $\underline{X}$. This desideratum is precisely satisfied by the static allocation defined in Equation 9, i.e.,
$\underline{X}=\sup\left\{X\mid X\leq\frac{B-\overline{\Delta}(X)}{\overline{N}}\right\},$
@@
where we let $\overline{N}=\mathbb{E}[N]+\textsc{Conf}_{0,T}^{N}$ for an appropriately defined confidence term $\textsc{Conf}_{0,T}^{N}$.
While $\underline{X}$ is budget-respecting with high probability, allocating a single budget-respecting quantity in each period may result in significant losses to efficiency. Herein lies the motivation behind the use of the “aggressive” allocation $\overline{X}=\underline{X}+L_{T}$: in periods in which the algorithm “feels” as though it is holding onto more units than it will need throughout the remainder of the horizon, it allocates aggressively. More specifically, in each period, our algorithm verifies whether there is enough budget remaining to accommodate $(i)$ the aggressive allocation $\overline{X}$ in the current period, $(ii)$ the lower guardrail $\underline{X}$ in all future periods, under high demand, and $(iii)$ aggressive future perishing. If enough budget remains, it allocates $\overline{X}$; otherwise, it defaults to $\underline{X}$. Formally, in any given period $t$, given a pessimistic forecast of future perishing $\overline{P}_{t}$, our threshold condition is as follows:
$B_{t}^{alg}\geq N_{t}\overline{X}+\overline{N}_{>t}\underline{X}+\overline{P}_{t},$
where we let $\overline{N}_{>t}=\mathbb{E}[N_{>t}]+\textsc{Conf}_{t,T}^{N}$ for an appropriately defined confidence term $\textsc{Conf}_{t,T}^{N}$.
Our main challenge lies in the construction of $\overline{P}_{t}$. In particular, computing a tight upper bound on future perishing poses significant challenges that do not exist in the setting without perishable resources. In this latter setting, it is easy to obtain tight upper bounds on future demand since demand uncertainty is exogenous, i.e., it is invariant to the decisions made by the algorithm. On the other hand, uncertainty around future perishing is endogenous: as discussed in Section 3, though the distribution around perishing times is fixed, how many — and which — items perish depends heavily on the rate at which items are being allocated, which itself depends on the rate at which items perish. We sidestep this challenge by turning to the “slow” consumption process introduced in Section 3. Namely, fix a round $t\in[T]$, and let $\tau_{b}(t\mid\underline{X},\sigma)$ denote the latest possible allocation time of unit $b$, starting from time $t$. (This latest possible allocation time is a time-dependent generalization of the latest possible allocation time defined in Section 3.2; we formally define it below.) In line with the intuition provided earlier, a necessary condition for $b$ to have spoiled after $t$ is for its perishing time $T_{b}$ to be such that $t\leq T_{b}<\min\{T,\tau_{b}(t\mid\underline{X},\sigma)\}$. We use this necessary condition to obtain a high-probability pessimistic estimate of the number of units that perish after period $t$, which we denote by $\overline{P}_{t}$. Crucially, $\overline{P}_{t}$ is independent of our algorithm’s decisions, which will allow for its tractable analysis. We formalize the construction of this key quantity below.
For all $t^{\prime}\geq t$, let $\underline{N}_{[t,t^{\prime}]}=\mathbb{E}\left[N_{[t,t^{\prime}]}\right]-\textsc{Conf}_{t,t^{\prime}}^{N}$ be a high-probability lower bound on the number of arrivals between $t$ and $t^{\prime}$, for an appropriately defined confidence term $\textsc{Conf}_{t,t^{\prime}}^{N}$. In line with the discussion in Section 3.2, with high probability, $\underline{N}_{<t}\underline{X}+\underline{N}_{[t,t^{\prime}]}\underline{X}$ corresponds to the smallest quantity of units that could have been consumed by $t^{\prime}$, since any units perishing before $t^{\prime}$ would only have increased consumption. Consider now any unit $b$ in the set of remaining items at the beginning of period $t$. Under the slow process defined above, if $\underline{N}_{<t}+\underline{N}_{[t,t^{\prime}]}\underline{X}\geq\sigma(b)$, then $b$ must have been allocated by $t^{\prime}\geq t$, according to the allocation schedule $\sigma$. We use this intuition to define $\tau_{b}(t\mid\underline{X},\sigma)$, i.e.,
${\tau}_{b}(t\mid\underline{X},\sigma)=\inf\left\{t^{\prime}\geq t:\underline{N}_{<t}\underline{X}+\underline{N}_{[t,t^{\prime}]}\underline{X}\geq\sigma(b)\right\}.$ (13)
Let $\overline{\mathcal{B}}_{t}$ denote the set of items remaining in period $t$ under this slow process. Via similar logic, $\overline{\mathcal{B}}_{t}=\{\sigma^{-1}(\lceil\underline{N}_{<t}\underline{X}\rceil),\ldots,\sigma^{-1}(B)\}$. Then, an upper bound on the expected number of items that perish from $t$ onwards is given by:
$\eta_{t}=\sum_{b\in\overline{\mathcal{B}}_{t}}\mathbb{P}\left(t\leq T_{b}<\min\{T,{\tau}_{b}(t\mid\underline{X},\sigma)\}\right).$ (14)
Here, we deliberately use $\overline{\mathcal{B}}_{t}$ — as opposed to $\mathcal{B}^{alg}_{t}$, which is endogenous to the allocation process — so that our pessimistic perishing estimate is completely exogenous to our algorithm’s decisions. This will be critical in our analysis, allowing us to sidestep a union bound over exponentially many subsets of remaining items.
With this machinery in hand, we define our high-probability pessimistic forecast $\overline{P}_{t}$ as follows:
$\begin{cases}\overline{P}_{t}&=\min\left\{\overline{P}_{t-1},\eta_{t}+\textsc{Conf}_{t}^{P}(\eta_{t})\right\}\quad\forall\ t\in[T],\\ \overline{P}_{0}&=B,\end{cases}$ (15)
for an appropriately defined confidence term $\textsc{Conf}_{t}^{P}(\cdot)$. Note that $\overline{P}_{1}=\overline{\Delta}(\underline{X})$ by construction, a fact we will use in the analysis of our algorithm.
@@
We present our algorithm, Perishing-Guardrail, in Algorithm 1. Recall, for $t\in[T]$, $\mathcal{B}^{alg}_{t}$ is used to denote the set of remaining resources at the beginning of time $t$, and $B_{t}^{alg}$ the quantity of remaining resources at the beginning of the period. Moreover, $\textsc{PUA}_{t}^{alg}$ denotes the quantity of unallocated items that perished at the end of round $t$.
@@
Input: Budget $B=B_{1}^{alg}$, allocation schedule $\sigma$, envy parameter $L_{T}$, arrival confidence terms $(\textsc{Conf}_{t,t^{\prime}}^{N})_{t,t^{\prime}\in\{0,\ldots,T\}}$, perishing confidence terms $(\textsc{Conf}_{t}^{P}(\cdot))_{t\in\{1,\ldots,T\}}$, and perishing inputs $(\eta_{t})_{t\in[T]}$ given by Equation 14
@@
Compute $\underline{X}=\sup\left\{X\mid X\leq\frac{B-\overline{\Delta}(X)}{\overline{N}}\right\}$ and set $\overline{X}=\underline{X}+L_{T}$.
@@
   Compute $\overline{P}_{t}=\min\left\{\overline{P}_{t-1},\eta_{t}+\textsc{Conf}_{t}^{P}(\eta_{t})\right\}$ // Compute ‘‘worst-case’’ future perishing
@@
 else if $B_{t}^{alg}-N_{t}\overline{X}\geq\underline{X}(\mathbb{E}\left[N_{>t}\right]+\textsc{Conf}_{t,T}^{N})+\overline{P}_{t}$ then// use upper guardrail